When the outstanding balance was $10,000 and a minimum payment of 3% is made, how many months does it take to pay off the balance?

approx. 92
approx. 198
approx. 244
approx. 290
Answer #1

FV= PV*(1+r)^n

here FV = future value which is zero

PV = initial balance = 10000

r= rate which is -3%

Hence

0=10000*(1-0.03)^n

0 = 10000*0.97^n

N= 290

At n= 290, Fv remaining = $1.46 which is negligible.